# Rigorous Test for Quantum Integrability and Nonintegrability

### v2

@@ IV Algorithm of integrability test
Next, we check whether <$\mathcal{B}^{(3)}_{<}$ is trivial (i.e., equal to $\mathcal{B}^{(2)}_{\leq}$) or not. This step is equivalent to finding solutions of the following equality:
@@
for $\hat{Q}_{[2],1}^{(2)}$ satisfying Eq. (50) and for $\hat{Q}_{[3],1}^{(2)}\in\mathcal{B}_{0}^{\otimes 2}$. If the above equality implies $\hat{Q}_{[2],1}^{(2)}=0$, we can conclude that this system is nonintegrable by Thm. 1. The above procedure consists of eigenvalue problems for a finite number of linear maps, which are independent of the system size.
As pointed out below Thm. 1, this algorithm can also be applied to a system of spin ladders and models with finite-range interactions. As discussed in Refs. [25, 26], the proof of nonintegrability of systems on hypercubic lattices with nearest-neighbor interactions can be reduced to that of ladder systems. Hence, we can indirectly deal with higher-dimensional systems as well.
Note that given additional pieces of information about the Hamiltonian, it may be possible to construct a less costly algorithm. For example, suppose that there exists a subspace $\mathcal{A}$ of $\mathcal{B}_{0}$ satisfying $\hat{H}_{1}^{(1)}\in\mathcal{A}$ and $\hat{H}_{1}^{(2)}\in\mathcal{A}^{\otimes 2}$. Then, by Eq. (50), we have $\hat{Q}_{[2],1}^{(2)}\in\mathcal{A}^{\otimes 2}$. Furthermore, a solution of Eq. (51) satisfies $\hat{Q}_{[3],1}^{(2)}\in\mathcal{A}_{2}^{\otimes 2}$, where $\mathcal{A}_{2}$ is defined as
$\mathcal{A}_{2}\coloneqq\operatorname{\mathrm{Span}}\{\hat{A}\hat{B}\mid\hat{A},\hat{B}\in\mathcal{A}\}\cap\mathcal{B}_{0}.$ (52)
This observation is useful to show the nonintegrability of general spin-$S$ systems. As a simple example, let us consider the spin-$S$ Heisenberg chain. In this case, $\mathcal{A}$ consists of a linear combination of spin operators $\hat{S}^{\alpha}$, and $\mathcal{A}_{2}$ is $8$-dimensional regardless of $S$ if $S\geq 1$. Therefore, we can show the nonintegrability of the Heisenberg chain for $S\geq 3/2$ as well as for $S=1$ [23, 24].
V Conclusion and outlook
@@
In this work, we have proven a sufficient condition using only 3-local quantities that guarantees nonintegrability, i.e., the absence of local conserved quantities. This result reveals a unified structure behind the existing proofs of nonintegrability, independent of the specific system, and provides a partial solution to the long-standing conjecture that the existence of 3-local conserved quantities is a necessary condition for integrability. Using the obtained theorem, we have also constructed a procedure for determining whether a given translationally invariant system is integrable or nonintegrable, independent of system size. This opens the way for numerical determination of nonintegrability with a rigorous basis.
@@
One straightforward extension of this work is a unified understanding of the proof of nonintegrability in systems that do not satisfy the assumption (4). Typical examples of such systems are those with Ising-type interactions, which are investigated in Refs. [17, 26]. In these proofs of nonintegrability, it may be necessary to consider up to $(k-1)$-local outputs to show the absence of $k$-local conserved quantities [17]. Thus, the structure of nonintegrability may be qualitatively different from the nonintegrability of the system satisfying the assumption (4).
@@
Another challenging direction is to discuss the existence or absence of structures other than local conserved quantities. In Sec. II.3, we have demonstrated that even in the MBL phase, there is no (strictly) local conserved quantity. In the context of distinguing the MBL phase and the ergodic phase where the ETH holds, a crucial question is whether the method of proofs of nonintegrability can be applied to quasi-local conserved quantities. As a related subject, we have discussed the absence of SGAs in Sec. III.4. If RSGAs could be analyzed, it would be possible to determine the presence or absence of quantum many-body scarring. These generalizations can lead to further classification of many-body systems beyond the presence or absence of local conserved quantities.
@@ Acknowledgements.
The author would like to thank Hosho Katsura, Mizuki Yamaguchi, Yuya Chiba, and Naoto Shiraishi for fruitful discuccions, and Masahito Ueda, Masaya Nakagawa and Hal Tasaki for their valuable comments on the manuscript. The author is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), World-Leading INnovative Graduate Study (WINGS) Program, The University of Tokyo.
@@ Appendix B Complete Proof of Theorem 2
Let $k\geq 4$. There is an isomorphism
$\mathcal{B}_{<}^{(k)}\ni\hat{Q}_{[k]}^{(k)}+\hat{Q}_{[k]}^{(k-1)}\mapsto\hat{Q}_{[k-1]}^{(k-1)}+\hat{Q}_{[k-1]}^{(k-2)}\in\mathcal{B}_{<}^{(k-1)},$ (54)
where $\hat{Q}_{[k]}^{(k)}$ and $\hat{Q}_{[k-1]}^{(k-1)}$ satisfy Eq. (24).
The proof of Prop. 8 follows the same procedure as for Prop. 6. First we construct a map from $\mathcal{B}_{<}^{(k)}$ to $\mathcal{B}_{<}^{(k-1)}$. For $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$ and $\hat{Q}_{[k]}^{(k-1)}\in\mathcal{B}^{(k-1)}$, the condition $\hat{Q}^{(k)}+\hat{Q^{\prime}}^{(k-1)}\in\mathcal{B}_{<}^{(k)}$ is equivalent to the following equalities:
for all $i\in\Lambda_{N}$, where $\hat{Q}_{[2]}^{(2)}$ is defined by Eq. (28). We denote the terms in the first and second brackets on the RHS of Eq. (55) as $\hat{F}^{(k)}_{i}$ and $\hat{G}^{(k)}_{i}$, respectively. Similarly to the case of $k=4$, we can decompose $\hat{F}^{(k)}_{i}$ and $\hat{G}^{(k)}_{i}$ in terms of operators $\hat{A}^{(k-1)}_{L},\hat{A}^{(k-1)}_{R},\hat{B}^{(k-1)}_{L}$, and $\hat{B}^{(k-1)}_{R}$. Specifically, we have the following lemma for $\hat{F}^{(k)}_{i}$.
@@
By combining the above two lemmas, we can show Prop. 8.
Proof of Prop. 8.
@@
$\hat{Q^{\prime}}^{(k-1)}_{L,i}$ $\coloneqq\hat{Q}^{(k-1)}_{[k],i}+\hat{A}^{(k-1)}_{L,i}+\hat{B}^{(k-1)}_{L,i},$ (60)
$\hat{Q^{\prime}}^{(k-1)}_{R,i}$ $\coloneqq\hat{Q}^{(k-1)}_{[k],i}-\hat{A}^{(k-1)}_{R,i}-\hat{B}^{(k-1)}_{R,i}.$ (61)
@@
$\hat{Q^{\prime}}^{(k-1)}_{L,i}-\hat{Q^{\prime}}^{(k-1)}_{R,i}=\hat{A}^{(k-1)}_{L,i}+\hat{A}^{(k-1)}_{R,i}+\hat{B}^{(k-1)}_{L,i}+\hat{B}^{(k-1)}_{R,i}.$ (62)
Furthermore, by using Eq. (55), we can find an operator $\hat{Q^{\prime}}^{(k-2)}_{i+1}\in\mathcal{B}_{0}^{\otimes(k-2)}$ satisfying
@@
for all $i\in\Lambda_{N}$, which is equivalent to $\hat{Q}_{[k-1]}^{(k-1)}+\hat{Q}_{[k-1]}^{(k-2)}\in\mathcal{B}_{<}^{(k-1)}$. A linear map from $\mathcal{B}_{<}^{(k)}$ to $\mathcal{B}_{<}^{(k-1)}$ is now constructed. By the same discussion as in the main text, this is in fact an isomorphism. ∎
The remaining task is to prove Lem. 9 and Lem. 10.
Proof of Lem. 9.
@@
$\hat{F}^{(k)}_{i,l}\coloneqq[\hat{Q}^{(k)}_{[k],i},\hat{H}^{(2)}_{i+l}].$ (66)
By using Lem. 5, we can decompose $\hat{F}^{(k)}_{i,l}$ for $31\leq l\leq k-$ 3 as

### v3

@@ V Algorithm of integrability test
Next, we check whether <$\mathcal{B}^{(3)}_{<}$ in the zero-momentum sector is trivial (i.e., equal to $\mathcal{B}^{(2,p=0)}_{\leq}$) or not. This step is equivalent to finding solutions of the following equation:
[Uncaptioned image] (78)
for $\hat{Q}_{[3],1}^{(2)}\in\mathcal{B}_{0}^{\otimes 2}$, where we assume that $\mathcal{B}_{\leq}^{(2,p=0)}=\mathbb{C}\hat{H}^{(2)}$. If Eq. (78) has no solution, we can conclude that this system has no $k$-local and translationally invariant conserved quantity for $23\leq k\leq N/$ 2. Furthermore, if $\mathcal{B}^{(2,p)}_{\leq}=\{0\}$ holds, or equivalently, Eq. (77) has no nontrivial solution for $p\neq 0$, then this system is nonintegrable. The above procedure involves eigenvalue problems for a finite number of linear maps, all of which are independent of system size. As pointed out below Thm. 1, this algorithm can also be applied to a system of spin ladders and models with finite-range interactions.
Note that if additional structural information about the Hamiltonian is available, a more efficient algorithm may be constructed. For example, suppose that there exists a subspace $\mathcal{A}$ of $\mathcal{B}_{0}$ satisfying $\hat{H}_{1}^{(2)}\in\mathcal{A}^{\otimes 2}$. Then, by the injectivity and Eq. (77), we have $\hat{Q}_{[2],1}^{(2)}\in\mathcal{A}^{\otimes 2}$. Furthermore, a solution of Eq. (78) satisfies $\hat{Q}_{[3],1}^{(2)}\in\mathcal{A}_{2}^{\otimes 2}$, where $\mathcal{A}_{2}$ is defined as
$\mathcal{A}_{2}\coloneqq\operatorname{\mathrm{Span}}\{\hat{A}\hat{B}-\operatorname{\mathrm{Tr}}(\hat{A}\hat{B})I,[\hat{A},\hat{H}_{1}^{(1)}]\mid\hat{A},\hat{B}\in\mathcal{A}\}.$ (79)
This observation is useful to show the nonintegrability of general spin-$S$ systems. As a simple example, let us consider the spin-$S$ Heisenberg chain. In this case, $\mathcal{A}$ consists of a linear combination of spin operators $\hat{S}^{\alpha}$, and $\mathcal{A}_{2}$ is eight-dimensional regardless of $S$ if $S\geq 1$. Therefore, we can show the nonintegrability of the XYZ chain for $S\geq 1$ as well as that of the Heisenberg chain for $S=1$ [23, 24], see Appx. C.
VI Nonintegrability of systems with internal degree of freedom
As pointed out in Ref. [25], the proof of nonintegrability for systems on hypercubic lattices and ladders with nearest-neighbor interactions can be reduced to that for ladder systems. In this section, we demonstrate that the problem of determining <$\mathcal{B}^{(k)}_{<}$ for systems with on-site interactions among internal degrees of freedom can be reduced to that of a ladder system. The Hamiltonian of such a system takes the following form:
$\hat{H}=\sum_{(\alpha,i)\in\mathcal{S}\times\Lambda_{N}}\hat{H}_{\alpha;i}^{(2)}+\sum_{i\in\Lambda_{N}}\sum_{\mathcal{S}_{0}\begin{subarray}{c}\subset\mathcal{S}\\ \neq\emptyset\end{subarray}}\hat{H}_{\mathcal{S}_{0};i}^{(1)},$ (80)
where $\hat{H}_{\alpha;i}^{(2)}\in\mathcal{B}_{0}^{\{(\alpha,i),(\alpha,i+1)\}}$ and $\hat{H}_{\mathcal{S}_{0};i}^{(1)}\in\mathcal{B}_{0}^{\mathcal{S}_{0}\times\{i\}}$ hold. Here, $\mathcal{S}$ denotes the set of labels for internal degree of freedom with $2\leq|\mathcal{S}|<\infty$. A typical example of such a model is the spin ladder system, where $|\mathcal{S}|=2$. Systems with nearest-neighbor interactions on the hypercubic lattice $(\Lambda_{N})^{D}$ correspond to the case $\mathcal{S}=(\Lambda_{N})^{D-1}$. In particular, systems with nearest-neighbor interactions on the honeycomb lattice can be regarded as a special case of that on the square lattice (see Fig. 2) and its Hamiltonian takes the form of Eq. (80).
Figure 2: The honeycomb lattice can be regarded as a subgraph of the square lattice. Therefore, a Hamiltonian with nearest-neighbor interactions on the honeycomb lattice takes the form of Eq. (80). We define locality along the direction indicated by the horizontal lines.
We denote the interaction and the potential on $\{\alpha\}\times\Lambda_{N}$ as $\hat{H}_{\alpha}$:
$\hat{H}_{\alpha}=\sum_{i\in\Lambda_{N}}\hat{H}_{\alpha;i}^{(2)}+\sum_{i\in\Lambda_{N}}\hat{H}_{\{\alpha\};i}^{(1)}.$ (81)
We assume the injectivity (4) of $\hat{H}_{\alpha}^{(2)}$ on the chain $\Lambda_{N}\times\{\alpha\}$ for all $\alpha\in\mathcal{S}$: for all $i\in\Lambda_{N}$,
$\hat{X}\in\mathcal{B}_{0}^{\{(\alpha,i+1)\}},\ [I\otimes\hat{X},\hat{H}^{(2)}_{\alpha;i}]=0$ $\Rightarrow\hat{X}=0$ (82)
$\mbox{and}\ \hat{Y}\in\mathcal{B}_{0}^{\{(\alpha,i)\}},\ [\hat{Y}\otimes I,\hat{H}^{(2)}_{\alpha;i}]=0$ $\Rightarrow\hat{Y}=0,$
which is equivalent to the injectivity of $\hat{H}^{(2)}$ on $\Lambda_{N}\times\mathcal{S}$.
Notably, the Hamiltonian (80) cannot satisfy the assumption (6). Precisely speaking, the following proposition holds.
Proposition 10.
For the Hamiltonian of the form given in Eq. (80) satisfying the injectivity (see Eq. (82)), the following holds:
$\mathcal{B}^{(2)}_{\leq}(\hat{H})=\bigoplus_{\alpha\in\mathcal{S}}\mathcal{B}^{(2)}_{\leq}(\hat{H}_{\alpha}).$ (83)
This relation is clear from the fact that $\hat{H}^{(2)}$ is regarded as the Hamiltonian of independent $|\mathcal{S}|$ chains. Since $\dim\mathcal{B}^{(2)}_{\leq}(\hat{H})=\sum_{\alpha}\dim\mathcal{B}^{(2)}_{\leq}(\hat{H}_{\alpha})\geq|\mathcal{S}|$ holds, the assumption (6) does not hold and we cannot show the nonintegrability of these systems using Thm. 1. However, we can provide another test for the triviality of <$\mathcal{B}^{(k)}_{<}(\hat{H})$. By combining Prop. 10 and Thm. 2, we can express $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}^{(k)}_{\leq}(\hat{H})$ as
$\hat{Q}_{[k]}^{(k)}$ $=\sum_{(\alpha,i)\in\mathcal{S}\times\Lambda_{N}}[\dots[[\hat{Q}_{[2],\alpha;i}^{(2)},\hat{H}_{\alpha;i+1}^{(2)}],\hat{H}_{\alpha;i+2}^{(2)}]\dots,\hat{H}_{\alpha;i+k-2}^{(2)}]$
$\eqqcolon\sum_{\alpha\in\mathcal{S}}\hat{Q}_{[k],\alpha}^{(k)}$ (84)
for some $\hat{Q}_{[2],\alpha}^{(2)}\in\mathcal{B}^{(2)}_{\leq}(\hat{H}_{\alpha})$. Then, we have the following theorem.
Theorem 11.
Consider the Hamiltonian of the form of Eq. (80) satisfying the injectivity (see Eq. (82)). For $23\leq k\leq N/$ 2 and $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}^{(k)}_{\leq}(\hat{H})$, if $\hat{Q}_{[k]}^{(k)}+\hat{Q}_{[k]}^{(k-1)}\in\mathcal{B}^{(k)}_{<}(\hat{H})$ holds for some $\hat{Q}_{[k]}^{(k-1)}\in\mathcal{B}^{(k-1)}$, then all of the following conditions hold.
  1. 1.
    For each $\alpha\in\mathcal{S}$, $\hat{Q}_{[k],\alpha}^{(k)}+\hat{Q}_{[k],\alpha}^{(k-1)}\in\mathcal{B}^{(k)}_{<}(\hat{H}_{\alpha})$ holds for some $(k-1)$-local operator $\hat{Q}_{[k],\alpha}^{(k-1)}$ on $\{\alpha\}\times\Lambda_{N}$.
  2. 2.
    For each $(\alpha,i)\in\mathcal{S}\times\Lambda_{N}$ and $\mathcal{S}_{0}\supsetneq\{\alpha\}$, the following holds:
    $[[\hat{Q}^{(2)}_{[2],\alpha;i},\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}],\hat{H}^{(2)}_{\alpha;i+1}]+[\hat{H}^{(2)}_{\alpha;i},[\hat{Q}^{(2)}_{[2],\alpha;i+1},\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}]]$
    $\hskip 42.67912pt=-(k-2)[[\hat{Q}^{(2)}_{[2],\alpha;i},\hat{H}^{(2)}_{\alpha;i+1}],\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}].$ (85)
  3. 3.
    For each $i\in\Lambda_{N}$, $\alpha\neq\beta\in\mathcal{S}$ and $\mathcal{S}_{0}\ni\alpha,\beta$, the following holds:
    $[[\hat{Q}^{(2)}_{[2],\alpha;i},\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}],\hat{H}^{(2)}_{\beta;i+1}]$
    $=[[\hat{H}^{(2)}_{\alpha;i},\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}],\hat{Q}^{(2)}_{[2],\beta;i+1}].$ (86)
Here, $\hat{Q}_{[k],\alpha}^{(k)}$ and $\hat{Q}_{[2],\alpha}^{(2)}$’s are defined through Eq. (84).
The first condition indicates that this system is no more integrable than a one-dimensional system $\hat{H}_{\alpha}$. In particular, if <$\mathcal{B}^{(k)}_{<}(\hat{H}_{\alpha})=\mathcal{B}^{(k-1)}_{\leq}(\hat{H}_{\alpha})$ holds for all $\alpha$, then $\hat{H}$ is nonintegrable. By combining with previous studies on the nonintegrability, This allows us to newly establish the nonintegrability of some systems, such as the quantum compass model on the cubic lattice.
The remaining two conditions are equations for 3-local quantities. The second involves a single chain with a “branch”, while the third describes the relation between two chains. From the theory of generalized eigenvalue problems [67], Eq. (85) either has solutions for all $k$, or only for finitely many (at most $\dim\mathcal{B}^{(2)}_{\leq}(\hat{H}_{\alpha})$) $k$. In particular, when $\mathcal{B}^{(2)}_{\leq}(\hat{H}_{\alpha})=\mathbb{C}\hat{H}^{(2)}_{\alpha}$ holds, Eq. (85) is equivalent, for all $k$, to the following equation:
$q_{\alpha}[[\hat{H}^{(2)}_{\alpha;i},\hat{H}^{(2)}_{\alpha;i+1}],\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}]=0,$ (87)
where we assume that $\hat{Q}^{(2)}_{[2],\alpha}=q_{\alpha}\hat{H}^{(2)}_{\alpha}$. This condition leads to the nonintegrability of models where $H_{a}$ is integrable, such as the XY(Z) model on the honeycomb lattice. In summary, the nonintegrability of the Hamiltonian in Eq. (80) reduces to a problem on at most two chains with a “branch”.
We conclude this section with the proof of Thm. 11.
Proof of Thm. 11.
We decompose $\hat{Q}_{[k]}^{(k-1)}$ as
$\hat{Q}_{[k]}^{(k-1)}=\sum_{\begin{subarray}{c}D\subset\mathcal{S}\times\Lambda_{N}\\ \operatorname{\mathrm{size}}{D}=k-1\end{subarray}}\hat{Q}_{[k],D}^{(k-1)},$ (88)
where the support of $\hat{Q}_{[k],D}^{(k-1)}$ is $D$. We define the following abbreviation for special forms of $D$ and the corresponding $\hat{Q}_{[k],D}^{(k-1)}$:
$D_{\alpha;i}^{(k)}$$\coloneqq\{\alpha\}\times\{i,\dots,i+k-1\},$ (89)
$\hat{Q}_{[k],\alpha;i}^{(k-1)}$ $\coloneqq\hat{Q}_{[k],D_{\alpha;i}^{(k-1)}}^{(k-1)},$ (90)
$D_{\alpha\mathcal{S}_{0}\beta;i}^{(k,l)}$ $\coloneqq D_{\alpha;i}^{(l)}\cup\mathcal{S}_{0}\times\{i+l\}\cup D_{\beta;i+l+1}^{(k-l-1)},$ (91)
$\hat{Q}_{[k],\alpha\mathcal{S}_{0}\beta;i}^{(k-1,l)}$ $\coloneqq\hat{Q}_{[k],D_{\alpha\mathcal{S}_{0}\beta;i}^{(k-1,l)}}^{(k-1)}.$ (92)
We set $D_{\alpha;i}^{(l)}=\emptyset$ for $l\leq 0$.
First, we show the first condition. By tracing out the commutator $[\hat{Q}_{[k]}^{(k)}+\hat{Q}_{[k]}^{(k-1)},\hat{H}]$ except for $\{\alpha\}\times\Lambda_{N}$, we obtain
$[\hat{Q}_{[k],\alpha}^{(k)}+\hat{Q}_{[k],\alpha}^{(k-1)},\hat{H}_{\alpha}]\in\tilde{\mathcal{B}}^{(k-1)}\cap\mathcal{B}^{\otimes(\{\alpha\}\times\Lambda_{N})},$
where $\mbox{where }\hat{Q}_{[k],\alpha}^{(k-1)}\coloneqq\sum_{i\in\Lambda_{N}}\hat{Q}_{[k],\alpha;i}^{(k-1)}.$ (93)
Here we use the assumption $\operatorname{\mathrm{len}}([\hat{Q}_{[k]}^{(k)}+\hat{Q}_{[k]}^{(k-1)},\hat{H}])\leq k-1$. It is clear that Eq. (93) is equivalent to $\hat{Q}_{[k],\alpha}^{(k)}+\hat{Q}_{[k],\alpha}^{(k-1)}\in\mathcal{B}^{k}_{<}(\hat{H}_{\alpha})$.
Next, we show Eq. (85) and Eq. (86). These equations are summarized as
$[[\hat{Q}^{(2)}_{[2],\alpha;i},\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}],\hat{H}^{(2)}_{\beta;i+1}]+[\hat{H}^{(2)}_{\alpha;i},[\hat{Q}^{(2)}_{[2],\beta;i+1},\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}]]$
$\hskip 36.98857pt=-\delta_{\alpha\beta}(k-2)[[\hat{Q}^{(2)}_{[2],\alpha;i},\hat{H}^{(2)}_{\alpha;i+1}],\hat{H}^{(1)}_{\mathcal{S}_{0};i+1}].$ (94)
By tracing out the commutator $[\hat{Q}_{[k]}^{(k)}+\hat{Q}_{[k]}^{(k-1)},\hat{H}]$ except for $D_{\alpha\mathcal{S}_{0}\beta;i}^{(k,l)}$ and focusing on $k$-local operators, we have
$[\hat{Q}_{[k],\alpha\mathcal{S}_{0}\beta;i}^{(k-1,l=0)},\hat{H}_{\beta;i+k-2}^{(2)}]+[\hat{Q}_{[k],\beta;i}^{(k)},\hat{H}_{\mathcal{S}_{0};i}^{(1)}]=0$ (95)
for $l=0$,
$[\hat{Q}_{[k],\alpha\mathcal{S}_{0}\beta;i}^{(k-1,l)},\hat{H}_{\beta;i+k-2}^{(2)}]+[\hat{Q}_{[k],\alpha\mathcal{S}_{0}\beta;i+1}^{(k-1,l-1)},\hat{H}_{\alpha;i}^{(2)}]$
$+\delta_{\alpha\beta}[\hat{Q}_{[k],\alpha;i}^{(k)},\hat{H}_{\mathcal{S}_{0};i+l}^{(1)}]=0$ (96)
for $21\leq l\leq k-$ 2 and
$[\hat{Q}_{[k],\alpha\mathcal{S}_{0}\beta;i+1}^{(k-1,l=k-2)},\hat{H}_{\alpha;i}^{(2)}]+[\hat{Q}_{[k],\alpha;i}^{(k)},\hat{H}_{\mathcal{S}_{0};i+k-1}^{(1)}]=0$ (97)
for $l=k-1$. These equations are translated by the string diagram as
@@ VI Nonintegrability of systems with internal degree of freedom
[Uncaptioned image] , (99)
[…]
Here, the labels under the output lines represent the degrees of freedom on which the operators act. By the injectivity, Eq. (100) yields
[…]
By using Eq. (99) iteratively, we have
[Uncaptioned image] . (102)
This equation and Eq. (98) give
[Uncaptioned image] , (103)
which is equivalent to Eq. (94). This concludes the proof. ∎
VII Conclusion and outlook
@@
In this work, we have proven a sufficient condition based solely on 3-local quantities that guarantees nonintegrability, i.e., the absence of local conserved quantities. This result reveals a unified structure underlying existing proofs of nonintegrability, independent of the specific system, and provides a partial resolution to the long-standing conjecture that the existence of 3-local conserved quantities is a necessary condition for integrability. Using the obtained theorem, we have also constructed a procedure for determining whether a given translationally invariant quantum spin chain is integrable or nonintegrable, independent of system size. This opens the way for numerical determination of nonintegrability with a rigorous basis.
@@
One straightforward extension of this work to develop a unified understanding of nonintegrability in systems that do not satisfy the assumption (4) or (6). Typical examples that do not satisfy the assumption (4) are systems with Ising-type interactions, which are investigated in Refs. [17, 26]. In these proofs of nonintegrability, it may be necessary to consider up to $(k-1)$-local outputs to show the absence of $k$-local conserved quantities [17]. Thus, the structure of nonintegrability may be qualitatively different from the nonintegrability of the system satisfying the assumption (4). The XY model is an example that satisfies the assumption (4) but fails to satisfy the assumption (6). Unlike the Ising-type interactions, the proof of the nonintegrability of this system [21, 22] is almost identical to that of systems satisfying both assumptions. Therefore, it is expected that our results can be extended to this case.
@@
Another promising direction is to explore the existence or absence of structures beyond integrability. In Sec. III.1, we have demonstrated that even in the MBL phase, there is no (strictly) local conserved quantity. In the context of distinguishing the MBL phase from the ergodic phase where the ETH holds, a crucial question is whether the methods used to prove nonintegrability can be extended to quasi-local conserved quantities. As a related topic, we have discussed the absence of SGAs in Sec. IV.4. If RSGAs could be analyzed, it may become possible to determine the presence or absence of quantum many-body scarring. These generalizations can lead to further classification of many-body systems beyond the presence or absence of local conserved quantities.
@@ Acknowledgements.
The author expresses sincere gratitude to Yuuya Chiba for carefully reviewing the manuscript and pointing out the gap in the proof. The author would like to thank Hosho Katsura, Mizuki Yamaguchi, and Naoto Shiraishi for fruitful discussions, and Masahito Ueda, Masaya Nakagawa, Hal Tasaki and Balázs Pozsgay for their valuable comments on the manuscript. The author is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), World-Leading INnovative Graduate Study (WINGS) Program, The University of Tokyo.
@@ Appendix B Complete Proof of Theorem 3
Let $k\geq 4$. There is a linear isomorphism
$\mathcal{B}_{<,b}^{(k)}\ni q\hat{Q}_{b}^{(k)}+\hat{Q}_{[k]}^{(k-1)}\mapsto q\hat{Q}_{b}^{(k-1)}+\hat{Q}_{[k-1]}^{(k-2)}\in\mathcal{B}_{<,b}^{(k-1)},$ (105)
where $q\in\mathbb{C}$ and $\hat{Q}_{b}^{(k)}$ is defined by Eq. (11).
The proof of Prop. 12 follows the same procedure as for Prop. 8. First we construct a map from $\mathcal{B}_{<,b}^{(k)}$ to $\mathcal{B}_{<,b}^{(k-1)}$. For $\hat{Q}_{[k]}^{(k-1)}\in\mathcal{B}^{(k-1)}$, the condition $q\hat{Q}_{b}^{(k)}+\hat{Q}_{[k]}^{(k-1)}\in\mathcal{B}_{<,b}^{(k)}$ is equivalent to the following equalities:
for all $i\in\Lambda_{N}$. We denote the terms in the first and second brackets on the RHS of Eq. (106) as $\hat{F}^{(k)}_{i}$ and $\hat{G}^{(k)}_{i}$, respectively. Similarly to the case of $k=4$, we can decompose $\hat{F}^{(k)}_{i}$ and $\hat{G}^{(k)}_{i}$ in terms of operators $\hat{A}^{(k-1)}_{L},\hat{A}^{(k-1)}_{R},\hat{B}^{(k-1)}_{L}$, and $\hat{B}^{(k-1)}_{R}$. Specifically, we have the following lemma for $\hat{F}^{(k)}_{i}$.
@@
By combining the above two lemmas, we can show Prop. 12.
Proof of Prop. 12.
@@
$\hat{Q^{\prime}}^{(k-1)}_{L,i}$ $\coloneqq\hat{Q}^{(k-1)}_{[k],i}+q\hat{A}^{(k-1)}_{L,i}+q\hat{B}^{(k-1)}_{L,i},$ (111)
$\hat{Q^{\prime}}^{(k-1)}_{R,i}$ $\coloneqq\hat{Q}^{(k-1)}_{[k],i}-q\hat{A}^{(k-1)}_{R,i}-q\hat{B}^{(k-1)}_{R,i}.$ (112)
@@
$\hat{Q^{\prime}}^{(k-1)}_{L,i}-\hat{Q^{\prime}}^{(k-1)}_{R,i}=q\left(\hat{A}^{(k-1)}_{L,i}+\hat{A}^{(k-1)}_{R,i}+\hat{B}^{(k-1)}_{L,i}+\hat{B}^{(k-1)}_{R,i}\right).$ (113)
Furthermore, by using Eq. (106), we can find an operator $\hat{Q^{\prime}}^{(k-2)}_{i+1}\in\mathcal{B}^{(k-2)}$ satisfying
@@
for all $i\in\Lambda_{N}$, which is equivalent to $q\hat{Q}_{b}^{(k-1)}+\hat{Q}_{[k-1]}^{(k-2)}\in\mathcal{B}_{<,b}^{(k-1)}$. A linear map from $\mathcal{B}_{<,b}^{(k)}$ to $\mathcal{B}_{<,b}^{(k-1)}$ is now constructed. By the same discussion as in the main text, this is in fact an isomorphism. ∎
The remaining task is to prove Lem. 13 and Lem. 14.
Proof of Lem. 13.
@@
$\hat{F}^{(k)}_{i,l}\coloneqq[\hat{Q}^{(k)}_{b,i},\hat{H}^{(2)}_{i+l}].$ (117)
By using Lem. 7, we can decompose $\hat{F}^{(k)}_{i,l}$ for $31\leq l\leq k-$ 3 as